OpenStudy (anonymous):

if n is a positive integer, generalize the binomial coefficient ( p ) for arbitrary n by setting: n ( p ) = ( p(p-1)....(p-n+1) ) / k! n put ( p ) = 1 0 a)show that sum ( p ) x^n converges for abs(x) <1 n b) show that (1+x)f'(x) -pf(x) =0 c) Deduce from b) that f(x) = 1+x)^p for abs x <1

OpenStudy (anonymous):

\[\binom pn=\frac{p(p-1)\cdots(p-n+1)}{n!}~~?\]

OpenStudy (anonymous):

The first question is to show that \[\large\sum_{n=0}^p\binom pnx^n~~\text{converges for }|x|<1\] which looks like a perfect candidate for the ratio test. The second question is a differential equation - a separable one. The third question could involve an application of the binomial theorem.

OpenStudy (anonymous):

@ssoleman got it? Or would you like some more in-depth help?

OpenStudy (anonymous):

@SithsAndGiggles sorry the site was down... its /k! not n! though...

OpenStudy (anonymous):

and do you just use the equation (p(p-1)...p-n+1 ) /k! to do the ratio test?

OpenStudy (anonymous):

I think the proff made a mistake and its n!

OpenStudy (anonymous):

Right, it has to be \(n\). There's no info about \(k\). First part: Ratio test: \[\begin{align*}\lim_{n\to\infty}\left|\frac{\dbinom p{n+1}x^{n+1}}{\dbinom pnx^n}\right|&=\lim_{n\to\infty}\left|\frac{\dfrac{p!}{(n+1)!(p-n-1)!}x^{n+1}}{\dfrac{p!}{n!(p-n)!}x^n}\right|\\\\ &=\lim_{n\to\infty}\left|\frac{1}{(n+1)(p-n-1)}x\right|\\\\ &=|x|\lim_{n\to\infty}\left|\frac{1}{(n+1)(p-n-1)}\right|\\\\ &=0\end{align*}\] Hmm, looks like the series converges for all \(x\)... Do I have the indices on the series right?

OpenStudy (anonymous):

yes thats right! I did that too for the first part

OpenStudy (anonymous):

now for b) we have to differentiate...but what do you mean by a seperable one?

OpenStudy (anonymous):

and i also dont understand how to show (1+x)f'(x) - pf(x) = 0 ... what is the -p(fx)??

OpenStudy (anonymous):

A separable differential equation is of the form \[A(x,y)\frac{dy}{dx}=B(x,y)\] which can be "separated" so that you can express the equation as a product of functions of different independent variables \(x\) and \(y\): \[a(x)b(y)\frac{dy}{dx}=\alpha(x)\beta(y)\] so that you can rewrite it as separate differentials: \[\frac{b(y)}{\beta(y)}~dy=\frac{\alpha(x)}{a(x)}~dx\] This allows you to integrate with respect to a single variable on both sides. You are given the differential equation \[(1+x)f'(x)-pf(x)=0\] Rewriting, you have \[\begin{align*} (1+x)f'(x)-pf(x)&=0\\ (1+x)f'(x)&=pf(x)\\ \frac{f'(x)}{f(x)}&=\frac{p}{1+x}\\ \frac{1}{f(x)}\frac{df(x)}{dx}&=\frac{p}{1+x}&\text{(another notation for derivative)}\\ \frac{1}{f(x)}~df(x)&=\frac{p}{1+x}~dx\\ \int\frac{1}{f(x)}~df(x)&=\int\frac{p}{1+x}~dx\\ \ln f(x)&=p\ln(1+x)+C\\ f(x)&=e^{p\ln(1+x)+C}\\ f(x)&=(e^{\ln(1+x)})^pe^C\\ f(x)&=C(1+x)^p\end{align*}\] Let me know if any of this doesn't make sense. Differential equations can be a lot to handle if this is your first time dealing with them. In solving the equation, you've established what you need to show for part (c).

OpenStudy (anonymous):

So what will part (c) look like?

OpenStudy (anonymous):

Part (c) seems to have at least two effective approaches. One is solving the differential equation above, and the other is an application of the binomial theorem. \[\large\sum_{n=0}^p\binom pna^nb^{p-n}=(a+b)^n\]\ Setting \(a=x\) and \(b=1\), you have \[\large\sum_{n=0}^p\binom pna^n1^{p-n}=\sum_{n=0}^p\binom pna^n=(x+1)^n\]\ Part (b) is a matter of deriving the differential equation, which isn't too difficult. It's a matter of differentiating the the given power series and algebraic rearrangements.